OpenStudy (anonymous):

Simplify. Assume that all nonnegative real numbers Square root of 12

OpenStudy (anonymous):

\[\sqrt{12}\]

OpenStudy (zzr0ck3r):

\[\sqrt{12}=\sqrt{4*3}=\sqrt{4}*\sqrt{3}\]